# The Type Ia Supernova 2021hem: A 2003fg-like Event in an Apparently Hostless Environment

### v1

@@ 5.1 A runaway progenitor from its host
Assuming SN 2021hem originated from a hyper-velocity runaway star (or binary system), we searched for known or visible galaxies in the vicinity of SN 2021hem as a plausible host. The closest galaxy candidate to SN 2021hem, as identified in the SIMBAD catalog, is an elliptical galaxy LEDA 1461683 ($\alpha=16^{h}21^{m}25^{s}.051$; $\delta=+14^{\circ}34^{\prime}45^{{}^{\prime\prime}}.60$; see Fig. 1), which has a redshift of $z=0.02873$. Although this value differs from the plausible galaxy group’s redshift of $\sim 0.035$ inferred in Sect. 3.4, such a discrepancy could be explained if the progenitor were a hyper-velocity star. A star with a line-of-sight velocity of $\sim 2000$ $\rm{\,km\,s^{-1}}$ would account for the higher observed redshift. Under this assumption, SN 2021hem would be located at a projected distance of $102$ kpc from the center of the galaxy. We noted that there is also a LINER-type Active Galaxy Nucleus, 2MASX J16212572+1431537, which is at a similar projected distance of 104 kpc and redshift of $z=0.02952$, making it an equally plausible host candidate as LEDA 1461683. However, we keep our analysis limited to LEDA 1461683, which is marginally closer to SN. As an interesting side note, considering the redshift and projected line of sight velocity of the star in this scenario, after accounting for reddening, luminosity distance, and K-correction, the absolute rest-frame magnitude of the SN would be $M_{B}=-19.45$ mag.
@@ 6 Conclusions
The Type Ia SN 2021hem was discovered within 48 hours of last non-detection in ZTF survey, in an apparently hostless environment. Spectral template matching yields a redshift of $z=0.0363\pm 0.0044$, which is also consistent with an overdensity of similar redshift galaxies within $\sim$1 Mpc radius. Its peak absolute $B$-band magnitude of $M_{B,max}=-19.96\pm 0.29$ mag places SN 2021hem at the bright end of the SN Ia luminosity distribution. Unlike normal or luminous 1991T-like SNe Ia, however, the $i$- and NIR-band light curves show no secondary maximum – a hallmark of low-luminosity 1991bg-like SNe Ia and some other peculiar subtypes. The high peak luminosity combined with the lack of a secondary maximum in the $i$- and NIR-band light curves suggests SN 2021hem is a 2003fg-like SN Ia candidate.
@@
Another intriguing aspect of SN 2021hem is the absence of a clearly associated host galaxy. 2003fg-like SNe Ia discoveries tend to show a preference towards low-mass dwarf host galaxies (e.g., 2011ApJ...733....3C; 2013MNRAS.432.3117T), however, these SNe have been found in a range of host-galaxy types, from low-mass dwarfs to massive spiral galaxies (2025A&A...694A..10D; 2025MNRAS.542.2752B). To constrain the host and progenitor of SN 2021hem, various possibilities have been thoroughly investigated. The nearest identified galaxy is located at a projected distance of 102 kpc, with a redshift of $z=0.02873$. Considering this as a plausible host, the progenitor could have been a hyper-velocity star (or binary system) ejected from it, traveling far off before the explosion. The transverse and radial velocity components of such a runaway hyper-velocity star can account for both the large offset from the host galaxy and the redshift difference between the host and that inferred for the SN. Assuming an explosion delay time of $100$ Myr for the SN, the estimated 3D unprojected velocity of this potential hyper-velocity progenitor is $\sim 2200$ km s-1.
@@
Taking into account all these factors, along with the limits we can place on any plausible undetected host galaxy, SN 2021hem remains one of the strongest candidates for a hostless supernova. Looking ahead, the combined capabilities of the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST), La Silla Quest Survey 4 (LS4; Miller2025), and The ATLAS Project (Tonry2018a) will greatly broaden the discovery space for such rare events.
@@ Appendix A Broad-band Photometry of SN 2021hem
Figure 12: The intrinsic $(\mbox{$B\!-\!V$})_{0}$ color curve of SN 2021hem corrected only for Milky-Way reddening is shown, and is compared with normal SN Ia 2011fe, and low-luminosity SN Ia 1991bg.
@@ Appendix C Probabilistic Multi-dimensional Clustering Analysis
As previously discussed in Section 4, SN 2021hem possesses many of the characteristics of 03fg-like SNe discussed in Ashall2021, (see their figure 18). As shown there, most of the 03fg-like objects are classified as “shallow-silicon” objects within the Branch scheme (Branch2006), with a couple objects classified as “core-normals”. Burrow2020 re-formulate the Branch Diagram using Gaussian Mixture Models, providing a method by which to estimate the probability any individual SN Ia belongs to each Branch group. Their clustering analysis also extended beyond the Branch Diagram phase space (i.e., pEW Si II $\lambda 5972$ vs pEW Si II $\lambda 6355)$ to include $v_{\rm Si\penalty 10000\ II}$ and $M_{B}$ in both three- and four-dimensional spaces. While the Branch groups are robust and tightly constrained even in higher-dimensional spaces, the groupings in these higher-dimensional spaces are by definition distinct from the original Branch groups.
Using the measurements from Ashall2021 and those of SN 2021hem determined in Sec. 4 above, we re-plot the traditional (e.g. two-dimensional) Branch Diagram in Fig. 14. Because group membership is determined probabilistically, we find that all 03fg-likes now belong to the “shallow-silicon” group, with a minimum membership probability of $p>0.79$ (see Table 5 for full probabilities). We find SN 2021hem is classified as a “core-normal” with $p=0.8526$, significantly offset from the rest of the 03fg-like sample. This is consistent with the placement of SN 2021hem in the $s_{BV}$ vs. $t^{i-B}_{max}$ space defined in Ashall2021 where it sits outside the region of 03fg-like SNe Ia.
Figure 14: The Branch Diagram including SN 2021hem (assuming $z=0.0363$), the sample of 03fg-like SNe Ia from Ashall2021 over-plotted with the Burrow2020 sample. We note that both CSS140501 and SN 2012dn have both moved from the “core-normal” classifications assigned in Ashall2021 to the “shallow-silicon” groups here. See Table 5 for the full probabilities, which correspond to the colorings of the overplotted points.
Table 5: Branch Group Membership Probabilities of 03fg-like SNe Ia
@@
Extending the above analysis to higher dimensional spaces by including $v_{\rm Si\penalty 10000\ II}$ and $M_{B}$ in the GMM clustering requires addressing the uncertainties in both quantities associated with the seeming hostless nature of SN 2021hem. We examine 3 cases: (1) assuming SN 2021hem lies at $z=0.0363$ as adopted in the main body of the paper; (2) the runaway progenitor case which assuming SN 2021hem lies within LEDA 1461683 at $z=0.02873$; and (3) assuming $z=0.0363$ but with larger uncertainties derived from propagating the redshift uncertainty as discussed in Sect. 3.4. We note that cases (1) and (2) consider only the measurement uncertainties associated with the Si II pEWs, $v_{\rm Si\penalty 10000\ II}$, and $M_{B}$. The results of this analysis are shown below in Fig. 15 and Table 6.
We find that that despite the overluminous nature of SN 2021hem, the inclusion of $M_{B}$ in both cases (1) and (3) changes the membership of SN 2021hem from the “core-normal”-equivalent group to the “cool”-equivalent group. The inclusion of $v_{\rm Si\penalty 10000\ II}$ causes the probabilities to shift 20% in the same direction without changing the group membership of SN 2021hemĠiven the well-established association between underluminous, rapidly-declining SNe Ia and the “Cool” Branch group (Branch2006; Burrow2020; Phillips2025); the placement of the overluminous, slowly-declining SN 2021hem in the “cool” group is likely a reflection of the abnormal, hostless nature of SN 2021hem, but may hint at our incomplete understanding of of 03fg-like SNe Ia. This placement likely arises due to: (1) a tight clustering of the “core-normal” objects along the $v_{\rm Si\penalty 10000\ II}$ and $M_{B}$ axes, (2) an upper-limit on pseudo-equivalent widths of the Si II lines for membership in the “shallow-silicon” group, and (3) the “broad-line” group being defined by high photospheric velocities to the point of being almost completely distinct from the other Branch groups in the higher dimensional spaces (Burrow2020).
In case (2), we find that SN 2021hem is associated with the “core-normal”-equivalent group at $p>0.77$ in all dimensionalities. SN 2021hem’s placement outside of the “shallow-silicon” (or equivalent) group in all cases and dimensionalities suggests that taken in suggesting that within this clustering scheme, its properties are significantly different than expected for a 03fg-like SNe Ia at $z=0.0363$. Instead, this method of analysis, taken in combination with the location of SN 2021hem in the Ashall2020 classification, is more consistent with SN 2021hem being a overluminous “core-normal” SN Ia arising from the hyper-velocity progenitor scenario discussed in Sect. 5.1. However, given both the diversity of 03fg-like SNe Ia behavior and the hostless nature of SN 2021hem, such results should be evaluated in combination with other analyses and remembering the large uncertainties on all measurements.
@@
Case 2: Runaway Progenitor ($z=0.02873$)
2D (Branch Diagram) 0.7773 ${4.817}\times 10^{-5}$ ${2.662}\times 10^{-6}$ 0.2227

### v2

@@ 5.1 A runaway progenitor from its host
Assuming SN 2021hem originated from a hyper-velocity runaway star (or binary system), we searched for known or visible galaxies in the vicinity of SN 2021hem as a plausible host. Two of the galaxies have been identified from SIMBAD (2000A&AS..143....9W) catalog as the closest and equidistant candidates from SN 2021hem (see Fig. 1). One is an elliptical galaxy LEDA 1461683 ($\alpha=16^{h}21^{m}25^{s}.051$; $\delta=+14^{\circ}34^{\prime}45\aas@@fstack{\prime\prime}60$) at $z=0.02873$, and a projected distance of $102\mbox{\penalty 10000\ kpc}$. The other is a LINER-type Active Galaxy Nucleus (AGN), 2MASX J16212572+1431537 ($\alpha=16^{h}21^{m}25^{s}.723$; $\delta=+14^{\circ}31^{\prime}53\aas@@fstack{\prime\prime}73$), which is at a similar projected distance of 104 kpc and redshift of $z=0.02952$. Although both of these galaxy redshifts differ from the plausible galaxy group’s redshift of $z\sim 0.035$ or the adopted SN redshift of 0.0363 as inferred in Sect. 3.4, such a discrepancy could be explained if the progenitor were a hyper-velocity star. A star with a line-of-sight velocity of $\sim 2000$ $\rm{\,km\,s^{-1}}$ would account for the higher observed redshift. As an interesting side note, considering the redshift and projected line-of-sight velocity of the star in this scenario, after accounting for reddening, luminosity distance, and K-correction, the absolute rest-frame magnitude of SN 2021hem would be $M_{B}=-19.45$ mag.
@@ 6 Conclusions
The Type Ia SN 2021hem was discovered within 48 hours of last non-detection in ZTF survey, in an apparently hostless environment. Spectral template matching yields a redshift of $z=0.0363\pm 0.0049$, which is also consistent with an overdensity of similar redshift galaxies within $\sim$1 Mpc radius. Its peak absolute $B$-band magnitude of $M_{B,max}=-19.96\pm 0.29$ mag places SN 2021hem at the bright end of the SN Ia luminosity distribution. Unlike normal or luminous 1991T-like SNe Ia, however, the $i$- and NIR-band light curves show no secondary maximum – a hallmark of low-luminosity 1991bg-like SNe Ia and some other peculiar subtypes. The high peak luminosity combined with the lack of a secondary maximum in the $i$- and NIR-band light curves suggests SN 2021hem is a 2003fg-like SN Ia candidate.
@@
Another intriguing aspect of SN 2021hem is the absence of a clearly associated host galaxy. 2003fg-like SNe Ia discoveries tend to show a preference towards low-mass dwarf host galaxies (e.g., 2011ApJ...733....3C; 2013MNRAS.432.3117T), however, these SNe have been found in a range of host-galaxy types, from low-mass dwarfs to massive spiral galaxies (2025A&A...694A..10D; 2025MNRAS.542.2752B). To constrain the host and progenitor of SN 2021hem, various possibilities have been thoroughly investigated. The nearest identified galaxy is an AGN, and is located at a projected distance of 104 kpc, with a redshift of $z=0.02952$. Considering this as a plausible host, the progenitor could have been a hyper-velocity star (or binary system) ejected from it, traveling far off before the explosion. An AGN host can facilitate the production of a hyper-velocity progenitor. The transverse and radial velocity components of such a runaway hyper-velocity star can account for both the large offset from the host galaxy and the redshift difference between the host and that inferred for the SN. Assuming an explosion delay time of $100$ Myr for the SN, the estimated 3D unprojected velocity of this potential hyper-velocity progenitor is $\sim 2200$ km s-1.
@@
Taking into account all these factors, along with the limits we can place on any plausible undetected host galaxy, SN 2021hem remains one of the strongest candidates for a hostless supernova. Looking ahead, the combined capabilities of the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; 2019ApJ...873.. [^1] I), La Silla Quest Survey 4 (LS4; Miller2025), and The ATLAS Project (Tonry2018a) will greatly broaden the discovery space for such rare events.
@@ Appendix A Broad-band Photometry of SN 2021hem
Figure 12: Intrinsic $(\mbox{$B\!-\!V$})_{0}$ color curve of SN 2021hem corrected only for Milky-Way reddening is shown, and is compared with normal SN Ia 2011fe, and low-luminosity SN Ia 1991bg.
@@ Appendix C Probabilistic Multi-dimensional Clustering Analysis
As previously discussed in Sect. 4, SN 2021hem possesses many of the characteristics of 2003fg-like SNe discussed in Ashall2021, (see their Fig. 18). As shown there, most of the 2003fg-like objects are classified as “shallow-silicon” objects within the Branch scheme (Branch2006), with a couple objects classified as “core-normals”. Burrow2020 re-formulate the Branch Diagram using Gaussian Mixture Models, providing a method by which to estimate the probability any individual SN Ia belongs to each Branch group. Their clustering analysis also extended beyond the Branch Diagram phase space (i.e., pEW Si ii $\lambda 5972$ vs. pEW Si ii $\lambda 6355)$ to include $v_{{\mathrm{Si\,\mathsc{ii}}}}$ and $M_{B}$ in both three- and four-dimensional spaces. While the Branch groups are robust and tightly constrained even in higher-dimensional spaces, the groupings in these higher-dimensional spaces are by definition distinct from the original Branch groups.
Using the measurements from Ashall2021 and those of SN 2021hem determined in Sect. 4.2, we re-plot the traditional (e.g. two-dimensional) Branch Diagram in Fig. 14. Because group membership is determined probabilistically, we find that all 2003fg-likes now belong to the “shallow-silicon” group, with a minimum membership probability of $p>0.79$ (see Table 5 for full probabilities). We find SN 2021hem is classified as a “core-normal” with $p=0.8526$, significantly offset from the rest of the 2003fg-like sample. This is consistent with the placement of SN 2021hem in the $s_{BV}$ vs. $t^{i-B}_{max}$ space defined in Ashall2021 where it sits outside the region of 2003fg-like SNe Ia.
Figure 14: Branch Diagram including SN 2021hem (assuming $z=0.0363$), the sample of 2003fg-like SNe Ia from Ashall2021 over-plotted with the Burrow2020 sample. We note that both CSS140501 and SN 2012dn have both moved from the “core-normal” classifications assigned in Ashall2021 to the “shallow-silicon” groups here. See Table 5 for the full probabilities, which correspond to the colorings of the overplotted points.
Table 5: Branch Group Membership Probabilities of 2003fg-like SNe Ia
@@
Extending the above analysis to higher dimensional spaces by including $v_{{\mathrm{Si\,\mathsc{ii}}}}$ and $M_{B}$ in the GMM clustering requires addressing the uncertainties in both quantities associated with the seeming hostless nature of SN 2021hem. We examine 3 cases: (1) assuming SN 2021hem lies at $z=0.0363$ as adopted in the main body of the paper; (2) the runaway progenitor case which assuming SN 2021hem lies within 2MASX J16212572+1431537 at $z=0.02952$; and (3) assuming $z=0.0363$ but with larger uncertainties derived from propagating the redshift uncertainty as discussed in Sect. 3.4. We note that cases (1) and (2) consider only the measurement uncertainties associated with the Si ii pEWs, $v_{\rm Si\penalty 10000\ II}$, and $M_{B}$. The results of this analysis are shown below in Fig. 15 and Table 6.
We find that that despite the overluminous nature of SN 2021hem, the inclusion of $M_{B}$ in both cases (1) and (3) changes the membership of SN 2021hem from the “core-normal”-equivalent group to the “cool”-equivalent group. The inclusion of $v_{\rm Si\penalty 10000\ II}$ causes the probabilities to shift 20% in the same direction without changing the group membership of SN 2021hem. Given the well-established association between underluminous, rapidly-declining SNe Ia and the “Cool” Branch group (Branch2006; Burrow2020; Phillips2025); the placement of the overluminous, slowly-declining SN 2021hem in the “cool” group is likely a reflection of the abnormal, hostless nature of SN 2021hem, but may hint at our incomplete understanding of of 2003fg-like SNe Ia. This placement likely arises due to: (1) a tight clustering of the “core-normal” objects along the $v_{\rm Si\penalty 10000\ II}$ and $M_{B}$ axes, (2) an upper-limit on pseudo-equivalent widths of the Si II lines for membership in the “shallow-silicon” group, and (3) the “broad-line” group being defined by high photospheric velocities to the point of being almost completely distinct from the other Branch groups in the higher dimensional spaces (Burrow2020).
In case (2), we find that SN 2021hem is associated with the “core-normal”-equivalent group at $p>0.8$ in all dimensionalities. SN 2021hem’s placement outside of the “shallow-silicon” (or equivalent) group in all cases and dimensionalities suggests that within this clustering scheme, its properties are significantly different than expected for a 2003fg-like SNe Ia at $z=0.0363$. Instead, this method of analysis, taken in combination with the location of SN 2021hem in the Ashall2020 classification, is more consistent with SN 2021hem being a overluminous “core-normal” SN Ia arising from the hyper-velocity progenitor scenario discussed in Sect. 5.1. However, given both the diversity of 2003fg-like SNe Ia behavior and the hostless nature of SN 2021hem, such results should be evaluated in combination with other analyses and remembering the large uncertainties on all measurements.
@@
Case 2: Runaway Progenitor ($z=0.02952$)
2D (Branch Diagram) 0.8018 ${8.450}\times 10^{-5}$ ${3.123}\times 10^{-6}$ 0.1981